OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $27.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? A. 600 B. 30 C. 150 D. 300

OpenStudy (anonymous):

Let c be the total cost and t be the time in minutes For Cell Plus: c=27+0.06t---(1 For Cell best: c=30+0.05t---(2 from ---(1 and ---(2 we have 27+0.06t =30+0.05t 0.06t-0.05t =30-27 0.01t= 3 t= 3/0.01 t= 300 Hence option D = 300 minutes is the correst Answer

OpenStudy (anonymous):

Hence, Trevor talk 300 minutes on both cell plans cost the same amount.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

U r most welcome